sin theta and cos theta are given. Find the exact value of the each of the four remaining trigonometric functions. Please explain.
sine theta = 2 sqrt5/5, cos theta = sqrt5/5

Answer by Alan3354(21605)

sin theta and cos theta are given. Find the exact value of the each of the four remaining trigonometric functions. Please explain.
sine theta = 2 sqrt5/5, cos theta = sqrt5/5
-----------
tan = sin/cos
cot = cos/sin
sec = 1/cos
csc = 1/sin
------